OpenStudy (anonymous):

The total length of two rivers is 13,385km. If the longer river were 205 km shorter, then it would be as long as the shorter river. Find the length of each river.

OpenStudy (anonymous):

x = short river length x+205 = longer river 2x + 205 = 13,385 2x = 13,180 x = 6,590 = shorter river longer one = 6,795

OpenStudy (anonymous):

You rock Nitin608!

OpenStudy (anonymous):

Let x be the length of the shorter river and y the length of the longer river. Then the total length = x + y = 13385 --(1) If the longer one is 205 km shorter, its length = y - 205 = x, the length of the shorter. ie x - y = - 205 --(2) Adding (1) and (2) 2x = 13180 Hence x = 13180/2 = 6590 kms Hence from (1) y = 13385 - x = 13385 - 6590 = 6795 kms Thus the lengths of the two rivers are 6590 kms and 6795 kms